OpenStudy (emmadrake):

x=24+xy; when x = -12 (please answer problem with steps)

jimthompson5910 (jim_thompson5910):

x=24+xy -12=24+(-12)y ... replace x with -12 -12 = 24 - 12y ------------- from here, we just isolate y. So we undo the stuff done to y to get it all by itself -12 = 24 - 12y -12-24 = 24 - 12y - 24 ... subtract 24 from both sides -36 = -12y -12y = -36 -12y/(-12) = -36/(-12) ... divide both sides by -12 y = 3

jimthompson5910 (jim_thompson5910):

In step 2, I subtracted 24 from both sides to undo the +24 In step 5, I undid the multiplication of -12 by dividing both sides by -12